OpenStudy (anonymous):

The top and bottom margins of a poster are 8 cm and the side margins are each 4 cm. If the area of printed material on the poster is fixed at 390 square centimeters, find the dimensions of the poster with the smallest area.

OpenStudy (anonymous):

for the width i got 21.96424004 but for some reason i keep getting the height wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

equation for Area: (x)(y)= 390 Separate for y: y= 390/x New area equation with border (x+ 8)(y+16) Multiply it out: (xy + 16x+8y+128 Now plug value for A initial in for xy , which is 390 390+ 16x +8y+128 Plug in y 390+16x+8(390/x)+128 = A Differentiate to optimize you'd get, then set to 0 to find relative minimum value 16-8(390/x^2)=0 x^2= 8*390 /16 x= 13.96 y=390/x = 390/13.96=27.928=27.93

OpenStudy (anonymous):

ok so the height is 21.9642004 and height = 27.93

OpenStudy (anonymous):

thank you!!